OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form.

OpenStudy (anonymous):

Step 1. Write down the generic form of Point-Slope y-y1=m(x-x1) (2,5); m=3 Step 2. Plug in values for x1 and y1, and your m (slope) y-5=3(x-2)

OpenStudy (anonymous):

thank you lots